OpenStudy (anonymous):

Hi, how could we solve this integration? ∫(dy/dx)^2 dx

OpenStudy (anonymous):

\[ \Large \int \frac{d^2y}{dx^2} \cdot dx \] ?

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=integrate+%28dy%2Fdx%29^2+with+respect+to+x i would try if i know there is solution.

OpenStudy (anonymous):

seems like an awkward notation to me in general. \[(f'(x))^2dx = f'(x)\cdot f'(x)dx \] Not that it doesn't exist, just haven't encountered it yet.

OpenStudy (turingtest):

I don't think you can simplify this as it is without more info about the function

OpenStudy (anonymous):

isn't that equal to \[\frac{dy}{dx}+C\]

OpenStudy (turingtest):

how would the exponent get lower @mukushla ?

OpenStudy (anonymous):

that's exactly what I would have derived next @experimentX q-:

OpenStudy (anonymous):

but I am really curious about the statement of the integral in general.

OpenStudy (anonymous):

lol....misread

OpenStudy (experimentx):

just expand the integrand wrt. Taylor series and integrate ... perhaps.

OpenStudy (anonymous):

i dont think they are equal @experimentX: \[\int\limits \left ( \frac{dy}{dx} \right )^2dx \neq \int\limits \left ( \frac{\partial y}{\partial x} \right )^2dx\] are they?

OpenStudy (turingtest):

it depends, like I said we need to know more about the function

OpenStudy (anonymous):

There is no closed solution for your integral.

OpenStudy (turingtest):

they could be equal I think, if the function is as such...

OpenStudy (turingtest):

That is what I think as well @eliassaab though I am having fun trying to integrate it by parts funny things are happening :)

OpenStudy (anonymous):

@TuringTest me too...parts

OpenStudy (experimentx):

yep ... so i scrupulously asked wolf.

OpenStudy (turingtest):

great minds think alike @mukushla ;)

OpenStudy (anonymous):

Do not waste your time. If Mathematica cannot do it, then it very probable that it is not feasible.

OpenStudy (anonymous):

its actually this one: \[\int\limits\limits v^2dt\]

OpenStudy (turingtest):

It may not be a waste of time trying something futile if you can learn exactly *why* it is futile I would say

OpenStudy (anonymous):

what is v?

OpenStudy (turingtest):

well what is \(v\) ?

OpenStudy (anonymous):

velocity

OpenStudy (anonymous):

\[v=\frac{dx}{dt}\] in physics

OpenStudy (turingtest):

then we are back to where we started

OpenStudy (experimentx):

x is some function of time right?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

should be if it's moving

OpenStudy (experimentx):

do you know that function?

OpenStudy (turingtest):

or even if it's not lol

OpenStudy (experimentx):

BTW why would you want to evaluate v^2?

OpenStudy (anonymous):

its the orginal problem: \[F=C\int\limits v^2dt\] where F is the force as a function of V and c is a constant

OpenStudy (experimentx):

can you state the original problem ... full, maybe ,,, we are missing something,

OpenStudy (anonymous):

oh! it would be so hard to explain!

OpenStudy (experimentx):

just try ... maybe we 'could' find the solution.

OpenStudy (anonymous):

so we could do something: \[dF=Cv^2dt \rightarrow \frac{dF}{dt}=m \frac{da}{dt}=m\frac{d^3x}{dt^3}=C \left( \frac{dx}{dt} \right)^2\]I think its not an integration anymore! it looks like a differential equation! can anyone solve this? \[\frac{d^3x}{dt^3}-\frac{C}{m} \left( \frac{dx}{dt} \right)^2=0\]or maybe i should ask it in an other question?

OpenStudy (experimentx):

non linear DE ...

OpenStudy (turingtest):

\[x'''-\frac Cm(x')^2=0\]let\[x'=y\]\[y''-\frac Cmy^2=0\implies y''=\frac Cmy^2\]I think we can solve that with undetermined coefficients

OpenStudy (experimentx):